OpenStudy (anonymous):

write a quadratic equation in standard form having the given solution -2,5/6

OpenStudy (luigi0210):

Set them equal to x, \(x=-2\) and \(x=\frac{5}{6}\). Now get them to equal 0.

OpenStudy (anonymous):

yes i did that

OpenStudy (anonymous):

(x+2)(x-5/6)

OpenStudy (anonymous):

=0

OpenStudy (luigi0210):

Don't you think it'd be easier to do \((5x-6\)?

OpenStudy (anonymous):

oooo i see

OpenStudy (anonymous):

(x+2)(5x-6)=0

OpenStudy (anonymous):

5x^2-6x+10x-12 thats what i got after doing foil

OpenStudy (anonymous):

my answer came up to be 5x^2+4x-12=0

OpenStudy (anonymous):

is that correct?

OpenStudy (luigi0210):

Yup, that's correct :)